The head of BPS Jakarta stated that the average consumption/capita of his regional household was USD 10,500,000 with a standard deviation of USD 2500,000. An economic research institute conducted a survey by selecting a random sample of 1000 households. The survey results show that the average consumption/capita is USD 12,000,000. Test it with an alpha of 2%, can you conclude that in fact the average consumption/capita is higher than that stated by the Head of BPS?


1
Expert's answer
2021-12-13T08:40:51-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu\leq10500000"

"H_1:\mu>10500000"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.02," and the critical value for a right-tailed test is "z_c =2.0537."

The rejection region for this left-tailed test is "R = \{z: z >2.0537\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{12000000−10500000}{2500000/\sqrt{1000}}\approx18.9737"

Since it is observed that "z = 18.9737>2.0537=z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z>18.9737)=0," and since "p=0<0.01=\alpha,"  it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that in fact the average consumption/capita is higher than that stated by the Head of BPS, at the "\alpha = 0.02." significance level.